Calculate the mass percent of carbon, nitrogen and oxygen in acetamide, C2H5NO


1
Expert's answer
2022-03-29T21:48:02-0400

Molar mass of C = 12g/mol

Molar mass of H = 1g/mol

Molar mass of N = 14g/mol

Molar mass of O = 16g/mol

Molar mass of C2H5NO = (12*2)+(1*5)+14+16 = 59g/mol


Mass percent of C = (2(12)/59)*100% = 40.7%

Mass percent of H = (1(5)/59)*100% = 8.5%

Mass percent of N = (14/59)*100 = 23.7%

Mass percent of O = (16/59)*100 = 27.1%


Note: make sure the percentages always add up to 100%

40.7+8.5+23.7+27.1= 100%